# Quantifying the Impact of the Dust Torque on the Migration of Low-mass Planets II: The Role of Pebble Accretion in Planet Growth within a Global Planet Formation Model

### v1

@@ Abstract
Although dust constitutes only about 1% of the mass of a protoplanetary disk, recent studies demonstrate that it can exert a significant torque on low- and intermediate-mass planetary cores. We compute and quantify for the first time the influence of the dust torque on the evolution of growing planetary embryos as they move in a protoplanetary disk while growing via gas and pebble accretion. Our global model evolves the gaseous disk via viscous accretion and X-ray photoevaporation, while accounting for dust growth and evolution including coagulation, drift, and fragmentation. Our research indicates that dust torque significantly influences planetary migration, particularly driving substantial outward migration for planets forming within the water ice-line. This effect occurs due to an increased dust-to-gas mass ratio in the inner disk, resulting from inward pebble drift from outer regions. In contrast, for planets initially located beyond the water ice-line, the dust torque mitigates inward migration but does not significantly alter their paths, as the dust-to-gas ratio diminishes rapidly due to rapid pebble drift and the brief timescales of planet formation in these areas. These findings underscore the pivotal role of dust torque in shaping the migration patterns of low- and intermediate-mass planets, especially when enhanced dust concentrations in the inner disk amplify its effects.
@@ 1 Introduction
Planet formation is a complex phenomenon that involves multiple physical processes occurring simultaneously, each playing a crucial role (see, e.g., Venturini et al., 2020c; Drążkowska et al., 2023; Mordasini & Burn, 2024, for recent reviews about planet formation). Several of these processes are closely interconnected. Notably, the growth of low- and intermediate-mass planets is heavily influenced by the evolution of both the gaseous and solid components in the protoplanetary disk, particularly within the framework of pebble accretion (e.g., Lambrechts et al., 2019; Liu et al., 2019; Ogihara & Hori, 2020; Venturini et al., 2020b, a; Drazkowska et al., 2021). The accretion of gas and solids is significantly affected by the orbital evolution of the growing planets, which is governed by their dynamical interactions with the disk in which they are embedded (e.g., Paardekooper et al., 2011; Lega et al., 2014; Benítez-Llambay et al., 2015; Jiménez & Masset, 2017).
@@
Most planet formation models including planet migration consider prescriptions that have been designed to account for the interactions between a planet and a gaseous disk (e.g., Ida & Lin, 2004; Alibert et al., 2005; Dittkrist et al., 2014; Guilera et al., 2017; Bitsch et al., 2019; Izidoro et al., 2021; Voelkel et al., 2022). Only a handful of models include the effects of thermal-disk diffusivity and heat released due to solid accretion onto the planet, both of which have an impact on planet migration in a gaseous disk (e.g., Guilera et al., 2019; Venturini et al., 2020a; Baumann & Bitsch, 2020; Guilera et al., 2021; Cornejo et al., 2023; Chrenko & Chametla, 2023).
@@
Even though the solid component is critical in most processes involved in planet formation and evolution, its role in planet migration has only been highlighted recently by Benítez-Llambay & Pessah (2018, hereafter BLP18), who showed that the solid component can generate a non-negligible torque even when contributing to the total disk mass at the standard 1% level (Drążkowska et al., 2023). This so-called “dust torque” is generally positive over a wide range of Stokes numbers and planet masses and has the potential to delay, and even reverse, the planet’s inward migration. This basic conclusion has also been confirmed by more recent works, which computed torque maps including the dynamical effects of pebble accretion onto the planet (Regály, 2020; Chrenko et al., 2024).
@@
In a previous paper (Guilera et al., 2023), based on torque measurements by BLP18, we built a framework to incorporate the dust torque on formation tracks models. We calculated the migration tracks of individual planets as they evolved embedded in various disk models based on the standard alpha-disk model. To gain insight into how different physical parameters influence the results, we examined the effects of independently varying the particles’ Stokes number, the dust-to-gas mass ratio, the stellar mass-accretion rate, and the disk viscosity. Through a series of “controlled experiments”, employing simple models for the evolution of the disk and the mass accretion onto the planet, we found that the dust torque can significantly impact the evolution of planetary embryos, leading to outward planet migration in many cases.
@@
The current paradigm for planet formation relies heavily on planetesimals embedded in dusty disks gaining mass through pebble accretion (see Ormel, 2024). This scenario involves combining elements that can lead to the impinging dusty flow to develop an asymmetric distribution around these accreting bodies, exerting a net torque on them (BLP18). In other words, planetesimals growing via pebble accretion are naturally subject to torques resulting from the flow of dust and pebbles passing by and accreting onto them. Currently, however, there are no global models of planet formation –based on pebble accretion– which include the orbital evolution of the growing planets as determined by the interaction between the planet, the gas, and the solid component in the protoplanetary disk.
@@ 2 Brief Description of the Model
Evolution of Solids – The evolution of the dust and pebbles is computed following Guilera et al. (2020), using a discrete size distribution with 200 size-bins between 1 $\mu$m and a maximum size. At each radial bin, the dust grows to a maximum size that is limited by coagulation, radial drift, and fragmentation –when viscosity becomes very low, fragmentation can be driven by differential drift (Birnstiel et al., 2012). We use the particle mass-weighted mean drift velocities and the mass-weighted mean Stokes numbers to compute the time evolution of the solid surface density solving an advection-diffusion equation (see, e.g., Dra̧zkowska et al., 2016).
@@
Accretion of Solids – We consider that planets grow by the concurrent accretion of pebbles and the surrounding gas as in Venturini et al. (2020b) and Guilera et al. (2021), and that the pebble accretion is halted when planets reach the pebble isolation mass (Lambrechts et al., 2014; Bitsch et al., 2018; Ataiee et al., 2018). As in Guilera et al. (2023), we adopt the conservative approach of allowing planets to grow until they reach $10~{}M_{\oplus}$ since, as shown in BLP18, the dust torque becomes inefficient beyond this mass range, when the planet opens a gap in the dust.
@@
Here, $\Gamma_{\text{g}}$ is the sum of the Lindblad and corotation torques exerted by the gas, for which we adopt the prescriptions given by Jiménez & Masset (2017). For the dust torque, $\Gamma_{\text{d}}$, we adopt the global models developed in Guilera et al. (2023), who provided torque maps resulting from a bi-dimensional linear interpolation using the planet mass and the mass-weighted mean Stokes numbers at the planet location based on the torque maps at a fiducial disk radius as computed by BLP18222In Appendix A, we compare the results of the model using the dust distribution instead the mass-weighted mean Stokes numbers to compute the dust torque..
@@ 3 Planet Formation Tracks
We adopt as our fiducial case study the disk model employed by Guilera et al. (2021). In summary, we consider a disk with a viscosity parameter $\alpha=10^{-4}$ with an initial mass of $M_{\rm disk}=0.05~{}M_{\odot}$ around a central solar-mass star, i.e., $M_{\star}=M_{\odot}$ with an initially constant dust-to-gas mass ratio $\epsilon(t=0~{}\text{yr})=0.01$. To understand the impact of the total disk mass and its metallicity, we consider in addition two complementary disk models: i) the fiducial case with a more massive disk, $M_{\rm disk}=0.1M_{\odot}$ and ii) the fiducial case with an initial metallicity $\epsilon(t=0~{}\text{yr})=0.02$. In what follows, we compute planet formation tracks for individual planets starting from an initial core of one Moon mass with a negligible envelope initially located at different disk radii, usually $r_{\rm p}(t=0~{}\text{yr})=\{0.5,1,2,3,5,8,10,12\}$ au.
@@ 3.1 Evolution of Fiducial Disk Model
Similar to the findings of Guilera et al. (2021), we observe that the mass-weighted mean Stokes numbers of particles are larger beyond the water ice-line, due to the higher threshold fragmentation velocities of ice-rich particles there. These Stokes numbers reach values between approximately $10^{-2}$ and $10^{-1}$ until the gap in the gas disk is formed. In contrast, within the water ice line, which is clearly distinguished as the boundary between the red and blue color scales, particles grow to smaller sizes, resulting in mass-weighted mean Stokes numbers in the range of $10^{-4}$ to $10^{-3}$. Following the formation of the gap, there is a significant reduction in the surface densities of gas and dust/pebbles, especially in the inner region of the disk where there is a lack of material supply. Consequently, dust and pebbles accumulate primarily in the pressure maximum created at the edge of the outer disk.
@@ 3.2 Planet Formation Tracks in Fiducial Disk Model
Figure 2: Planet formation tracks (planet total mass vs. semi-major axis) for the fiducial disk. Planets inside 2 au begins their formations inside the water ice-line. The dashed lines in panel b) represent the planet core masses.
@@
Figure 3: Individual torque components and the total torque on the planet, as a function of the planet mass for the planet initially located at 0.5 au from the fiducial run.
@@
The reason why the planets initially located at 0.5 and 1 au ultimately have lower masses when considering the dust torque is illustrated in Figure 5. In the absence of the dust torque (red lines), planets grow and continuously migrate through a region where the dust-to-gas ratio is at its highest; significantly boosting the pebble accretion rates. However, when the dust torque is taken into account (black lines), both planets initially migrate outward to a region with a reduced dust-to-gas mass ratio, resulting in less efficient growth.
Figure 4: Planet mass versus time for the planets initially located at 0.5 and 1 au (solid and dashed lines). The black/red lines distinguish whether the dust torque is considered/ignored.
@@ 3.3 Planet Formation Tracks in Massive Disks
In Figure 6, we present the time evolution maps of the particle mass-weighted mean Stokes numbers of the dust distribution (top panel) and the dust-to-gas mass ratio (bottom panel) for a massive disk with an initial mass of $0.1~{}M_{\odot}$ solar masses. The time evolution in each of these panels closely resembles their counterparts in Figure 1 because of a compensatory effect. Although dust can grow to slightly larger sizes — limited by fragmentation in the inner disk, which depends on gas surface density, and by drift in the outer disk, which depends on dust surface density (e.g., Birnstiel et al., 2012; Dra̧zkowska et al., 2016; Guilera et al., 2021; Birnstiel, 2023) — the Stokes numbers are directly proportional to dust sizes but inversely proportional to gas surface density, which is higher in this scenario. As a result, there are no significant differences in Stokes numbers across the disk between the moderate-mass and massive disk cases. Notably, the main differences compared to the fiducial case include the opening of the gas gap at approximately 3.5 Myr and the initial location of the water ice-line at around 3 au.
In Figure 7, we display the planet formation tracks for the case of the massive disk. Once again, for planets located beyond the water ice-line (i.e., those initially positioned beyond 3 au), the dust torque is not significant despite the higher Stokes numbers of pebbles in that region. This is primarily due to the rapid formation timescales of these planets and the swift decline of the dust-to-gas ratio in that region of the disk (see Guilera et al., 2023). Similar to the previous section, for planets situated inside the water ice-line, the dust torque plays a crucial role. In all scenarios, the dust torque dominates the total torque during the early stages, causing planets—especially those starting at 0.5 au and 1 au—to migrate outwards. The dust torque significantly alters the planet formation tracks in all cases, resulting in planets that are less massive compared to scenarios where the dust torque is not considered. As demonstrated by Guilera et al. (2023), the dust torque is more effective due to the increased dust-to-gas ratio (see Figure 6), even with the relatively low Stokes numbers inside the water ice-line. Furthermore, as observed in the previous section, the final planet masses are reduced because the dust torque drives planets to migrate outward into regions with lower dust-to-gas ratios, which subsequently decreases the pebble accretion rates.
Figure 6: Time evolution maps of the particle mass-weighted mean Stokes numbers (top) and the dust-to-gas mass ratio $\epsilon=\Sigma_{\text{d}}/\Sigma_{\text{g}}$ (bottom) for the case of a massive disk of $M_{\rm disk}=0.1M_{\odot}$.

### v2

@@ Abstract
Although dust constitutes only about 1% of the mass in a protoplanetary disk, recent studies reveal its substantial impact on the torques experienced by low- and intermediate-mass planetary cores. In this study, we present the first comprehensive analysis of the dust torque’s influence on the evolution of growing planetary embryos as they migrate through a protoplanetary disk and undergo gas and pebble accretion. Our global model incorporates viscous accretion and X-ray photoevaporation effects on the gaseous disk while also accounting for the dynamic processes of dust growth and evolution, including coagulation, drift, and fragmentation. Our findings demonstrate that dust torque significantly affects planetary migration patterns, particularly facilitating prominent outward migration for planets forming within the water ice line. This outward thrust arises from an enhanced dust-to-gas mass ratio in the inner disk, driven by the inward drift of pebbles from the outer regions. Conversely, for planets that originate beyond the water ice line, while the dust torque attenuates inward migration, it does not substantially alter their overall migration trajectories. This is attributed to the rapid reduction in dust-to-gas mass ratio, resulting from swift pebble drift and the short formation timescales prevalent in that region. Overall, our findings highlight the critical role of dust torque in shaping the migration of low- and intermediate-mass planets, particularly in conditions where increased dust concentrations amplify its effects. These insights have significant implications for understanding the formation timescales, mass distributions, and compositional characteristics of emerging planetary systems.
@@ 1 Introduction
Planet formation is a complex phenomenon that involves multiple physical processes occurring simultaneously, each playing a crucial role (see, e.g., Venturini et al., 2020c; Drazkowska et al., 2023; Mordasini & Burn, 2024, for recent reviews about planet formation). Several of these processes are closely interconnected. Notably, the growth of low- and intermediate-mass planets is heavily influenced by the evolution of both the gaseous and solid components in the protoplanetary disk, particularly within the framework of pebble accretion (e.g., Lambrechts et al., 2019; Liu et al., 2019; Ogihara & Hori, 2020; Venturini et al., 2020b, a; Drazkowska et al., 2021). The accretion of gas and solids is significantly affected by the orbital evolution of the growing planets, which is governed by their dynamical interactions with the disk in which they are embedded (e.g., Paardekooper et al., 2011; Lega et al., 2014; Benítez-Llambay et al., 2015; Jiménez & Masset, 2017).
@@
Most planet formation models including planet migration consider prescriptions that have been designed to account for the interactions between a planet and a gaseous disk (e.g., Ida & Lin, 2004; Alibert et al., 2005; Dittkrist et al., 2014; Guilera et al., 2017; Bitsch et al., 2019; Izidoro et al., 2021; Voelkel et al., 2022). Only a handful of models include the effects of thermal-disk diffusivity and heat released due to solid accretion onto the planet, both of which have an impact on planet migration in a gaseous disk (e.g., Guilera et al., 2019; Venturini et al., 2020a; Baumann & Bitsch, 2020; Guilera et al., 2021; Cornejo et al., 2023a, b; Chrenko & Chametla, 2023).
@@
Even though the solid component is critical in most processes involved in planet formation and evolution, its role in planet migration has only been highlighted recently by Benítez-Llambay & Pessah (2018, hereafter BLP18), who showed that the solid component can generate a non-negligible torque even when contributing to the total disk mass at the standard 1% level (Drazkowska et al., 2023). This so-called “dust torque” is generally positive over a wide range of Stokes numbers and planet masses and has the potential to delay, and even reverse, the planet’s inward migration. This basic conclusion has also been confirmed by more recent works, which computed torque maps including the dynamical effects of pebble accretion onto the planet (Regály, 2020; Chrenko et al., 2024) and also through linear analysis (Hou & Yu, 2024, 2025).
@@
In a recent paper, we developed a framework to integrate dust torque into formation track models, utilizing torque measurements from BLP18 (Guilera et al., 2023). We calculated the migration tracks of individual planets as they evolved embedded in various disk models based on the standard alpha-disk model. To gain insight into how different physical parameters influence the results, we examined the effects of independently varying the particles’ Stokes number, the dust-to-gas mass ratio, the stellar mass-accretion rate, and the disk viscosity. Through a series of simulations, employing simple models for the evolution of the disk and the mass accretion onto the planet, we found that the dust torque can significantly impact the evolution of planetary embryos, leading to outward planet migration in many cases.
@@
The current paradigm for planet formation relies heavily on planetesimals embedded in dusty disks gaining mass through pebble accretion (see Ormel, 2024, for a recent review). This scenario involves combining elements that can lead to the impinging dusty flow to develop an asymmetric distribution around these accreting bodies, exerting a net torque on them (BLP18). In other words, planetesimals growing via pebble accretion are naturally subject to torques resulting from the flow of dust and pebbles passing by and accreting onto them. Currently, however, there are no global models of planet formation –based on pebble accretion– which include the orbital evolution of the growing planets as determined by the interaction between the planet, the gas, and the solid component in the protoplanetary disk.
@@ 2 Brief Description of the Model
Evolution of Solids – The evolution of the dust and pebbles is computed following Guilera et al. (2020), using a discrete size distribution with 200 size-bins between 1 $\mu$m and a maximum size. At each radial bin, the dust grows to a maximum size that is limited by coagulation, radial drift, and fragmentation –when viscosity becomes very low, fragmentation can be driven by differential drift (Birnstiel et al., 2012). We use the particle mass-weighted mean drift velocities and the mass-weighted mean Stokes numbers to compute the time evolution of the solid surface density solving an advection-diffusion equation (see, e.g., Drazkowska et al., 2016).
@@
Accretion of Solids – As in Venturini et al. (2020b) and Guilera et al. (2021), we consider that planets grow by the concurrent accretion of surrounding pebbles and gas. Pebble accretion is computed in the 2D and 3D regime considering the pebble scale height (Lambrechts et al., 2014; Brasser et al., 2017). The accretion of pebbles is halted when the planets reach the pebble isolation mass (Lambrechts et al., 2014; Bitsch et al., 2018; Ataiee et al., 2018). Gas accretion is considered in both the attached and detached phases, following the approach of Venturini et al. (2020b). During the attached phase, when the planet’s envelope remains smoothly connected to the surrounding gaseous disk, the accretion process is modeled by solving the one-dimensional, spherically symmetric structure equations for the envelope, up until the planet reaches the pebble isolation mass. At this point, solid accretion ceases, and the core becomes critical. In this regime, where solid accretion is no longer occurring, we employ the prescription from Ikoma et al. (2000). As the planet surpasses the pebble isolation mass, its gas accretion rate may exceed the supply regulated by the disk’s viscosity. This leads to the detachment of the planet from the disk, while accreting gas at a rate dictated by viscosity, i.e., $3\pi\nu\Sigma_{\rm gas}$. Additionally, in the detached phase, gas accretion can be further suppressed if the planet opens a gap in the disk (Tanigawa & Ikoma, 2007). As in Guilera et al. (2023), we adopt the conservative approach of allowing planets to grow until they reach $10~{}M_{\oplus}$ since, as shown in BLP18, the dust torque becomes inefficient beyond this mass range, when the planet opens a gap in the dust.
@@
Here, $\Gamma_{\text{g}}$ is the sum of the Lindblad and corotation torques exerted by the gas, for which we adopt the prescriptions given by Jiménez & Masset (2017). For the dust torque, denoted as $\Gamma_{\text{d}}$, we utilize the global models created by Guilera et al. (2023). They generated smooth torque maps through a two-dimensional linear interpolation based on the planet’s mass and the mass-weighted mean Stokes number [^2] In Appendix A, we compare the results of the model using the dust distribution instead the mass-weighted mean Stokes number to compute the dust torque. at the planet’s location, using the torque maps calculated by BLP18 for a reference disk radius. The detailed computations and numerical implementations of the migration of planets in PLANETALP can be found in Guilera et al. (2019).
@@ 3 Planet Formation Tracks
We adopt as our fiducial case study the disk model employed by Guilera et al. (2021). Our disk has an initial gas surface density given by (Andrews et al., 2010)
$\Sigma_{\text{gas}}=\Sigma_{\text{gas}}^{0}\left(\dfrac{R}{R_{c}}\right)\,e^{-(R/R_{c})},$ (2)
$\Sigma_{\text{gas}}^{0}$ being a normalization constant that depends on the disk mass, and, $R_{c}=39$ au, is the characteristic radius of the disk. We also consider a viscosity parameter $\alpha=10^{-4}$ with an initial mass of $M_{\rm disk}=0.05~{}M_{\odot}$ around a central solar-mass star, i.e., $M_{\star}=M_{\odot}$ with an initially constant dust-to-gas mass ratio $\epsilon(t=0~{}\text{yr})=0.01$. To understand the impact of the total disk mass and its metallicity, we consider in addition two complementary disk models: i) the fiducial case with a more massive disk, $M_{\rm disk}=0.1M_{\odot}$ and ii) the fiducial case with an initial metallicity $\epsilon(t=0~{}\text{yr})=0.02$  [^3] In all cases we check the gravitional stablity of the disk by computing the Toomre-paremeter (Toomre, 1964) at every timestep and every radial bin (see Ronco et al., 2017, for details). All disks presented in this work are gravitationally stable.. In what follows, we compute planet formation tracks for individual planets starting from an initial core of one Moon mass and with a negligible envelope, initially located at different disk radii. We usually consider $r_{\rm p}(t=0~{}\text{yr})=\{0.5,1,2,3,5,8,10,12\}$ au.
@@ 3.1 Evolution of Fiducial Disk Model
As the threshold fragmentation velocity of the ice-rich particles is larger beyond the water ice-line, they reach larger sizes and Stokes numbers with respect to the silicate particles inside the water-ice line. This leads to mass-weighted mean Stokes numbers that are larger beyond the water ice-line. These Stokes numbers reach values between approximately $10^{-2}$ and $10^{-1}$ until the gap in the gas disk is formed. In contrast, within the water ice line (clearly distinguishable as the boundary between the red and blue color scales in Figure 1), particles grow to smaller sizes, resulting in mass-weighted mean Stokes numbers in the range of $10^{-4}$ to $10^{-3}$. After the gap forms, there is a marked decrease in the surface densities of gas and dust/pebbles, particularly in the inner region of the disk where material supply is limited. Consequently, dust and pebbles accumulate primarily in the pressure maximum created by the internal photoevaporation at the edge of the outer disk where the aerodynamical drag force vanishes.
@@ 3.2 Planet Formation Tracks in Fiducial Disk Model
Figure 2: Planet formation tracks (planet total mass vs. semi-major axis) for the fiducial disk. Planets inside 2 au begin their formations inside the water ice-line. The dashed lines in panel b) represent the masses of planet’s core.
@@
Figure 3: Individual torque components and the total torque on the planet, as a function of the planet mass for the planet initially located at 0.5 au in the fiducial run.
@@
Figure 5 illustrates why the planets initially positioned at 0.5 and 1 au end up with lower masses when accounting for the effects of dust torque. In the absence of the dust torque (red lines), planets grow and continuously migrate through a region where the dust-to-gas ratio is highest; significantly boosting the pebble accretion rates. However, when the dust torque is taken into account (black lines), both planets initially migrate outward to a region with a reduced dust-to-gas mass ratio, resulting in less efficient growth.
Figure 4: Planet mass versus time for the planets initially located at 0.5 and 1 au (solid and dashed lines). The black/red lines indicate whether the dust torque is considered/ignored.
@@ 3.3 Planet Formation Tracks in Massive Disks
In Figure 6, we present the time evolution maps of the particle mass-weighted mean Stokes numbers of the dust distribution (top panel) and the dust-to-gas mass ratio (bottom panel) for a massive disk with an initial mass of $0.1~{}M_{\odot}$ solar masses. The time evolution in each of these panels closely resembles their counterparts in Figure 1 because of a compensatory effect. Although dust can grow to slightly larger sizes — limited by fragmentation in the inner disk, which depends on gas surface density, and by drift in the outer disk, which depends on dust surface density (e.g., Birnstiel et al., 2012; Drazkowska et al., 2016; Guilera et al., 2021; Birnstiel, 2024) — the Stokes numbers are directly proportional to dust sizes but inversely proportional to gas surface density, which is higher in this scenario. As a result, there are no significant differences in mass-weighted mean Stokes numbers across the disk between the moderate-mass and massive disk cases. Notably, the main differences compared to the fiducial case include the opening of the gas gap at approximately 3.5 Myr and the initial location of the water ice-line at around 3 au.
In Figure 7, we display the planet formation tracks for the case of the massive disk. Once again, for planets located beyond the water ice-line (i.e., those initially positioned beyond 3 au), the dust torque is not significant in spite of the higher Stokes numbers of pebbles in that region. This is primarily due to the rapid formation timescales of these planets and the swift decline of the dust-to-gas ratio in that region of the disk (see Guilera et al., 2023). Similar to the fiducial disk model, for planets situated inside the water ice-line the dust torque plays a crucial role. In all scenarios, the dust torque dominates the total torque during the early stages, causing planets (especially those starting at 0.5 au and 1 au) to migrate outwards. The dust torque significantly alters the planet formation tracks in all cases, resulting in planets that are less massive compared to scenarios where the dust torque is not considered. As demonstrated by Guilera et al. (2023), the dust torque is more effective due to the increased dust-to-gas ratio (see Figure 6), even with the relatively low Stokes numbers inside the water ice-line. Additionally, as noted in the previous section, the final planet masses decrease because the dust torque pushes planets to migrate outward into areas with lower dust-to-gas ratios, which in turn reduces the rates of pebble accretion.
Figure 6: Time evolution maps of the particle mass-weighted mean Stokes numbers (top) and the dust-to-gas mass ratio $\epsilon=\Sigma_{\text{d}}/\Sigma_{\text{g}}$ (bottom) for the case of a massive disk with $M_{\rm disk}=0.1M_{\odot}$.